A number is 4 times the sum of its digits.The units digit is two more than the tens digit.Find the number.
Answer by ankor@dixie-net.com(22740)

Let x = the 10's digit
Let y = the units
then
10x+y = "a number"
:
A number is 4 times the sum of its digits.
10x + y = 4(x + y)
10x + y = 4x + 4y
10x - 4x = 4y - y
6x = 3y
Simply, divide by 3
2x = y
;
The units digit is two more than the tens digit.
y = x + 2
Replace y with 2x
2x = x + 2
2x - x = 2
x = 2
then
y = 4
:
24 is the number
:
:
See if the works in the statement:
"A number is 4 times the sum of its digits."
24 = 4(2 + 4)



Find the number.